PLEASE, HELP ME WITH THIS PROBLEM! I'VE BEEN TRYING SINCE YESTERDAY AND I COULDN'T FINISH YET =( THANKS.
In 8h, two campers rowed 15min down a river and then rowed back their campsite. The rate of the river's current was 1mph. Find the rate at which that campers row in calm water.

Answer by scott8148(6628)

assuming min is mi. __ t=d/r __ let x=rate in calm water

15/(r+1) + 15/(r-1) = 8 __ multiplying by (r+1)(r-1) __ 15r-15 + 15r+15 = 8r^2-8

8r^2-30r-8=0 __ dividing by 2 __ 4r^2-15r-4=0

factoring __ (4r+1)(r-4)=0

4r+1=0 __ r=-1/4 __ negative value not realistic

r-4=0 __ r=4 mph